OpenStudy (anonymous):

Can anyone help me with rationalizing denominators? I do not understand the process

OpenStudy (anonymous):

\[1/\sqrt{96}\]

OpenStudy (anonymous):

Multiply the numerator and the denominator by the denominator if it is a square root. Then simplify the expression.

OpenStudy (anonymous):

so, multiply 96 by 96 and get 1/9216?

OpenStudy (anonymous):

No, multiply the whole fraction by \[\sqrt{96}\]

OpenStudy (anonymous):

You will get \[\sqrt{96}/96\]